OpenStudy (anonymous):

Isolate the variable, sketch a reference triangle, find all the solutions within the given domain 0≤x≤2π sin²x+cos2x-cosx=0

OpenStudy (anonymous):

@zepdrix please help

OpenStudy (tkhunny):

Have you considered: 1) \(\sin^{2}(x) = 1 - \cos^{2}(x)\) 2) \(cos(2x) = 2\cos^{2}(x) - 1\)

OpenStudy (anonymous):

@tkhunny thanks, i got it now